# Training-Conditional Coverage Bounds under Covariate Shift

### v3

@@ 3.2 Full conformal and jackknife+
Assumption 2
The model is bi-Lipschitz (Lipeomorphism) in parameters, i.e.,
$\kappa_{1}\big{\|}\beta-\beta^{\prime}\big{\|}_{\infty}\leq\big{\|}\mu_{\beta}-\mu_{\beta^{\prime}}\big{\|}_{\infty}\leq\kappa_{2}\big{\|}\beta-\beta^{\prime}\big{\|}_{\infty},$
with $\kappa_{1}>0$ and $\kappa_{2}<\infty$.
@@
${F^{\prime(n)}}<L_{n}$ and ${F^{\prime(n)}_{Q}}<L_{Q,n}$.
@@
Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Assume $Q\ll P$ and $dQ/dP\leq B$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Assume $Q\ll P$ and $dQ/dP\leq B$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@ 4 Discussion
We have presented training-conditional coverage guarantees for conformal prediction methods under two covariate shift regimes: (1) $\|dQ/dP\|_{\infty}<\infty$ and (2) $\|dQ/dP\|_{P,2}<\infty$. The split conformal method has been analyzed for the black-box training algorithm while the full conformal and jackknife+ have been analyzed under three assumptions. Although Assumptions 1 and 2 have been verified only for the ridge regression in this paper, we conjecture that they are satisfied by a truncated version of the general RKHS regressions which we leave for future research. Truncated and sketched versions of the RKHS models have been extensively studied in the previous literature from the computational efficiency perspective (see Amini (2021); Williams and Seeger (2000); Zhang et al. (2013b); Alaoui and Mahoney (2015); Cortes et al. (2010) and references therein). The results in this paper quantify the training sample size (or calibration sample size for split conformal) needed for the coverage under covariate shift.
The $(m,n)$-stability parameters were recently introduced in Liang and Barber (2023) and used to compute training-conditional coverage bounds for inflated full conformal and jackknife+ prediction intervals under exchangeability. Unlike uniform stability which is a distribution-free property of a training process, $(m,n)$-stability depends on both the training algorithm and the distributions of the data as follows
$\psi_{m,n}^{\text{out}}$ $=\operatorname{\mathbb{E}}_{{\bf D}_{n+m}}|\hat{\mu}_{{\bf D}_{n}}(X_{\text{test}})-\hat{\mu}_{{\bf D}_{n+m}}(X_{\text{test}})|,$ (9)
$\psi_{m,n}^{\text{in}}$ $=\operatorname{\mathbb{E}}_{{\bf D}_{n+m}}|\hat{\mu}_{{\bf D}_{n}}(X_{1})-\hat{\mu}_{{\bf D}_{n+m}}(X_{1})|,$ (10)
with $X_{\text{test}}\perp\!\!\!\perp{\bf D}_{n+m}$ and ${\bf D}_{n+m}=((X_{1},Y_{1}),...,(X_{n+m},Y_{n+m}))$. Although weaker than uniform stability, these parameters are yet not well-understood in a practical sense. To elaborate on the difference between this approach and uniform stability, we evaluate their resulting training-conditional bounds for the ridge regression under exchangeability.
@@
[…]
@@
Letting $m^{-1/2}=(m/n)^{1/3}$ to balance the two terms $\sqrt{\frac{\log(1/\delta)}{\min(m,n)}}$ and $3]{{mc_{n-1}/\gamma}}$, we get $m=n^{2/5}$. By plugging $m=n^{2/5}$ into (12), we have
@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
Let $F_{Q}(x)=\operatorname{\mathbb{P}}_{Z\sim Q}(s(Z)\leq x)$ where $s(\cdot)$ denotes some fixed score function. Define,
@@
Assume $Q\ll P$ and $dQ/dP\leq B$. Then, for all $\delta>0$
$\operatorname{\mathbb{P}}\Bigg{(}\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta+2C\sqrt{\frac{B}{n}}\Bigg{)}\leq 4e^{-n\delta^{2}/(2B^{2})}.$
@@
$=\sup_{x\in\mathbb{R}}\left|\frac{1}{n}\sum_{i\in[n]}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\{s(Z_{i})\leq x\}-F_{Q}(x)\right|+\left|1-\frac{1}{n}\sum_{i\in[n]}\frac{dQ}{dP}(Z_{i})\right|$ (15)
Regarding the first term we have,
@@
$=\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|$
@@
[…]
according to the Holder’s inequality. Hence, $\|dQ/dP\|_{P,2}\leq\sqrt{B}$ and byVan Der Vaart and Wellner [1996, Theorem 2.14.2] we have
[…]
with some universal constant $C>0$, and $N_{[]}\left(\epsilon\|dQ/dP\|_{P,2},\mathcal{F},L^{2}(P)\right)$ denotes the bracketting number for function class =$\mathcal{F}=\{f_{x}:x\in\mathbb{R}\}$. It remains to compute an upper bound for $N_{[]}\left(\epsilon,\mathcal{F},L^{2}(P)\right)$. Let $t_{0}=-\infty$ and define,
[…]
@@
where $10\leq i\leq k-$ 1. Clearly, $\mathcal{F}\subset\cup_{i=0}^{k}b_{i}$ with $b_{k}=\{f\}$. Brackets $b_{0},\ldots,b_{k-1}$ have size
[…]
according to (17) and bracket $b_{k}$ has size 00. Hence, we have $k+1$ brackets in total with size smaller than $\epsilon$. By (18), we have $k-1\leq(\|dQ/dP\|_{P,2}/\epsilon)^{2}$ which implies that $N_{[]}\left(\epsilon,\mathcal{F},L^{2}(P)\right)\leq k+1\leq 2+(\|dQ/dP\|_{P,2}/\epsilon)^{2}$. Plugging this into (18), we get
@@
since
@@
$\Bigg{|}\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P)\right|-\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P^{(i)}_{n}-P)\right|\Bigg{|}$ $\leq\sup_{x\in\mathbb{R}}\left|\int f_{x}~{}d(P_{n}-P^{(i)}_{n})\right|$
@@
Therefore, by McDiarmid’s inequality McDiarmid et al. [1989] we get
@@
Back to (15), for the second term Hoeffding’s inequality implies
@@
$\operatorname{\mathbb{P}}\Bigg{(}\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta\,\Bigg{)}\leq\frac{6\,CK}{\delta\sqrt{n}}+\frac{4(K^{2}-1)}{n\delta^{2}}.$
@@
[…]
@@
Proof: To prove this lemma, we compute the upper bound for the first term of (15) differently. According to Yukich [1985, Theorem 2.1] and Lafferty et al. [2008, Theorem 7.86] we get
@@
for $\epsilon\leq 2/3$, where $N_{[]}(\epsilon/8,\mathcal{F},L^{1}(P))$ denotes the bracketting number for function class =$\mathcal{F}=\{f_{x}:x\in\mathbb{R}\}$. Similar to the proof of Lemma 1, it can be shown that $N_{[]}(\epsilon,\mathcal{F},L^{1}(P))\leq 2+1/\epsilon$. Therefore,
@@
Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
@@
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}\right)^{2}(Z_{1})\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{1})\leq n\right\}\right)},$
@@
$\left|F_{Q}(x)-F_{Q|\mathcal{X}_{B}}(x)\right|$ $=\left|\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\right)-\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)\right|$
@@
$\leq\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)\Bigg{|}\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big{|}\frac{dQ}{dP}(Z_{\text{test}})>B\right)$
@@
$\leq\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right)$
@@
$\operatorname{\mathbb{P}}\bigg{(}\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|$ $\leq\delta+C^{\prime}\sqrt{B/n_{B}}\,\Big{|}\,n_{B}\bigg{)}\geq 1-4e^{-n_{B}\delta^{2}/(2B^{2})}$

### v4

@@ 3.2 Full conformal and jackknife+
Assumption 2 (Bi-Lipschitz continuity)
The map $\beta\mapsto\mu_{\beta}$ is bi-Lipschitz with respect to the $\infty$-norms, i.e.,
$\kappa_{1}\,\|\beta-\beta^{\prime}\|_{\infty}\;\leq\;\|\mu_{\beta}-\mu_{\beta^{\prime}}\|_{\infty}\;\leq\;\kappa_{2}\,\|\beta-\beta^{\prime}\|_{\infty},$
for some constants $0<\kappa_{1}\leq\kappa_{2}<\infty$.
@@
${F^{\prime(n)}}<L_{n}$ and ${F^{\prime(n)}_{Q}}<L_{Q,n}$ where $F^{\prime(n)}$ and $F^{\prime(n)}_{Q}$} denote the derivative of $F^{(n)}$ and $F^{(n)}_{Q}$}, respectively.
@@
Define $P_{e}^{\text{J+}}({\bf D}_{n}):=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{J+}}_{\alpha}(X_{\text{test}})\Big|{\bf D}_{n}\right)$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Now we present our result on jackknife+ under covariate shift when $dQ/dP\leq B$.
@@
Assume that $Q$ is absolutely continuous with respect to $P$ ($Q\ll P$) and $dQ/dP\leq B$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Define $P_{e}^{\text{CV+}}({\bf D}_{n}):=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{CV+}}_{\alpha}(X_{\text{test}})\Big|{\bf D}_{n}\right)$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Define $P_{e}^{\text{full}}({\bf D}_{n}):=\operatorname{\mathbb{P}}\left(Y_{\text{test}}\not\in\hat{C}^{\text{full}}_{\alpha}(X_{\text{test}})\Big|{\bf D}_{n}\right)$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@
Assume that $Q$ is absolutely continuous with respect to $P$ ($Q\ll P$) and $dQ/dP\leq B$. Under Assumptions 13, for all $\epsilon,\delta>0$, it holds that
@@ 4 Discussion
We have presented training-conditional coverage guarantees for weighted conformal prediction methods under covariate shift. The split conformal method has been analyzed for the black-box training algorithm while the full conformal and jackknife+ have been analyzed under three assumptions. Although Assumptions 1 and 2 have been verified only for the ridge regression in this paper, we conjecture that they are satisfied by a truncated version of the general RKHS regressions which we leave for future research. Truncated and sketched versions of the RKHS models have been extensively studied in the previous literature from the computational efficiency perspective (see Amini (2021); Williams and Seeger (2000); Zhang et al. (2013b); Alaoui and Mahoney (2015); Cortes et al. (2010) and references therein). The results in this paper quantify the training sample size (or calibration sample size for split conformal) needed for the coverage under covariate shift.
Estimation of the likehood ratio. A natural question here is how to include the error arising from estimating the likelihood ratios in the analysis. Define
$\hat{w}_{i}:=\frac{V_{i}}{V}:=\frac{\frac{dQ}{dP}(Z_{i})}{\sum_{i\in[n]}\frac{dQ}{dP}(Z_{i})},\hskip 50.00008pt\breve{w}_{i}:=\frac{\hat{V}_{i}}{\hat{V}}:=\frac{\widehat{\frac{dQ}{dP}}(Z_{i})}{\sum_{i\in[n]}\widehat{\frac{dQ}{dP}}(Z_{i})}$
and let $s(\cdot)$ denote a fixed score function. In this case, one needs to include the following term to the weighted DKW inequalities derived in this paper,
$\sup_{t\in\mathbb{R}}\left|\sum_{i\in[n]}(\breve{w}_{i}-\hat{w}_{i})\boldsymbol{1}\{s(Z_{i})\leq t\}\right|$ $\leq\sum_{i\in[n]}|\breve{w}_{i}-\hat{w}_{i}|$
$=\sum_{i\in[n]}\left|\frac{\hat{V}_{i}}{\hat{V}}-\frac{V_{i}}{V}\right|$
$\leq\frac{\sum_{i\in[n]}\left(\left|\hat{V}_{i}V-V_{i}V\right|+\left|V_{i}V-V_{i}\hat{V}\right|\right)}{V\hat{V}}$
$=\frac{V\sum_{i\in[n]}\left|\hat{V}_{i}-V_{i}\right|+\left|V-\hat{V}\right|\left(\sum_{i\in[n]}V_{i}\right)}{V\hat{V}}$
$=\frac{\sum_{i\in[n]}\left|\hat{V}_{i}-V_{i}\right|+\left|V-\hat{V}\right|}{\hat{V}}\leq 2\frac{\sum_{i\in[n]}\left|\hat{V}_{i}-V_{i}\right|}{\hat{V}}$,
where the concentration properties of this term around zero depends on the estimator of $dQ/dP(Z_{i})$.
Comparison with the $(m,n)$-stability. The $(m,n)$-stability parameters were recently introduced in Liang and Barber (2023) and used to compute training-conditional coverage bounds for inflated full conformal and jackknife+ prediction intervals under exchangeability. Unlike uniform stability which is a distribution-free property of a training process, $(m,n)$-stability depends on both the training algorithm and the distributions of the data as follows
$\psi_{m,n}^{\text{out}}$ $:=\operatorname{\mathbb{E}}|\hat{\mu}_{{\bf D}_{n}}(X_{\text{test}})-\hat{\mu}_{{\bf D}_{n+m}}(X_{\text{test}})|,$ (9)
$\psi_{m,n}^{\text{in}}$ $:=\operatorname{\mathbb{E}}|\hat{\mu}_{{\bf D}_{n}}(X_{1})-\hat{\mu}_{{\bf D}_{n+m}}(X_{1})|,$ (10)
with $X_{\text{test}}\perp\!\!\!\perp{\bf D}_{n+m}$, ${\bf D}_{n+m}=((X_{1},Y_{1}),...,(X_{n+m},Y_{n+m}))$, and $\hat{\mu}_{{\bf D}_{n}}$ denotes the predictor function obtained by training on ${\bf D}_{n}$. Although weaker than uniform stability, these parameters are yet not well-understood in a practical sense. To elaborate on the difference between this approach and uniform stability, we evaluate their resulting training-conditional bounds for the ridge regression under exchangeability.
@@
$\operatorname{\mathbb{P}}\Bigg(P_{e}^{\text{J+},\gamma}({\bf D}_{n})>\alpha+3\sqrt{\frac{\log(1/\delta)}{\min(m,n)}}+$ $2\sqrt[3]{\frac{\psi_{m,{n-1}}^{\text{out}}}{\gamma}}\Bigg)\leq 3\delta+\sqrt[3]{\frac{\psi_{m,{n-1}}^{\text{out}}}{\gamma}},$ (11)
@@
Letting $m^{-1/2}=(m/n)^{1/3}$ to balance the two terms $\sqrt{\frac{\log(1/\delta)}{\min(m,n)}}$ and $\sqrt[3]{{mc_{n-1}/\gamma}}$, we get $m=n^{2/5}$. By plugging $m=n^{2/5}$ into (12), we have
@@
5 Conclusion
In this work, we have studied the training-conditional coverage bounds of full conformal, jackknife+, and CV+ prediction regions from a uniform stability perspective, which is well-understood for convexly regularized empirical risk minimization over reproducing kernel Hilbert spaces. We have derived new bounds via a concentration argument for the (estimated) predictor function. In the case of ridge regression, we have used the uniform stability parameter to derive a bound for $(m,n)$-stability and compare the resulting bounds from Liang and Barber (2023) to the bounds established in this paper. We have observed that our rates are faster in sample size but dependent on the dimension of the problem. Even though our work is theoretical in nature, it can potentially shed light on understanding a much broader downstream-task setups in reconstructive self-supervised learning (SSL), where the existing literature focuses on either linear regression or ridge regression (e.g., Lee et al. (2021); Teng et al. (2022); Du and Xiang (2024)). For split conformal, our result allows for flexible downstream setups in SSL well beyond simple regressions, while for jackknife+ it would be interesting to reveal the interplay between uniform stability and excess risk analysis in SSL. Another worthwhile direction is from the robustness perspective in conformal prediction and detection problems, including adversarial and label noise settings (e.g., Gendler et al. (2021); Einbinder et al. (2024); Zhang et al. (2025)). For instance, one can study the robustness of training conditional guarantees under adversarial attacks during inference time.
Acknowledgment
This work was supported in part by the National Science Foundation under Grant CCF-2611415.
@@ Appendix A Technical Lemmas: Weighted DKW Inequalities
A.1 Background
In this section, we first briefly introduce the notion of bracketing number and four foundamental results, which we will be applying in our proofs.
Bracketing Numbers. A measure of complexity for a class of functions is the bracketing number. For a pair of functions $f$ and $g$, the bracket $[f,g]$ is defined as follows,
$f,g]=\{h:f\leq h\leq g\}.$
A set of brackets $f_{1},g_{1}],\ldots,[f_{n},g_{n}]$ is a $\epsilon$-$L^{q}(P)$ bracketing of function class $\mathcal{F}$ if $\mathcal{F}\subseteq\bigcup_{1\leq i\leq n}[f_{i},g_{i}]$ and $(\int|f_{i}-g_{i}|^{q}dP)^{1/q}\leq\epsilon$ for all $i\in[n]$. The bracketing number $N_{[]}\left(\epsilon,\mathcal{F},L^{q}(P)\right)$ is the size of the smallest $\epsilon$-$L^{q}(P)$ bracketing of $\mathcal{F}$.
Theorem 7 (Van Der Vaart and Wellner [1996](Theorem 2.14.2))
Let $\mathcal{F}$ be a class of measurable functions with a measurable envelope $F$. Then, for some universal constant $C>0$,
$\operatorname{\mathbb{E}}\left(\sup_{f\in\mathbb{\mathcal{F}}}\left|\sqrt{n}\int f~d(P_{n}-P)\right|\right)\leq C\|F\|_{P,2}\int_{0}^{1}\sqrt{1+\log N_{[]}\left(\epsilon\|F\|_{P,2},\mathcal{F},L^{2}(P)\right)}~d\epsilon.$ (14)
Theorem 8 (Lafferty et al. [2008](Theorem 7.86))
Let $\mathcal{F}$ be a class of measurable functions with $A=\sup_{f\in\mathcal{F}}\|f\|_{P,1}$ and $B=\sup_{f\in\mathcal{F}}\|f\|_{\infty}$. Then, for $\epsilon<2A/3$,
$\operatorname{\mathbb{P}}\left(\sup_{f\in\mathbb{\mathcal{F}}}\left|\int f~d(P_{n}-P)\right|>\epsilon\right)\leq 4N_{[]}\left(\epsilon/8,\mathcal{F},L^{1}(P)\right)\exp{\left(-\frac{96n\epsilon^{2}}{76AB}\right)}.$ (15)
Theorem 9 (Durrett [2019] (Theorem 2.2.11))
For each $n$ let $X_{n,k}$, $1\leq k\leq n$ be independent. Let $b_{n}>0$ with $b_{n}\to\infty$, and let $\overline{X}_{n,k}=X_{n,k}\boldsymbol{1}\{|X_{n,k}|\leq b_{n}\}$. Suppose that, as $n\to\infty$, (i) $\sum_{k=1}^{n}\operatorname{\mathbb{P}}(|X_{n,k}|>b_{n})\to 0$, and (ii) $b_{n}^{-2}\sum_{k=1}^{n}\operatorname{\mathbb{E}}\overline{X}_{n,k}^{2}\to 0$. If we let $S_{n}=X_{1,n}+\ldots+X_{n,n}$ and put $a_{n}=\sum_{k=1}^{n}\operatorname{\mathbb{E}}\overline{X}_{n,k}$, then $(S_{n}-a_{n})/b_{n}\to 0$ in probability.
Proof: Fix $\varepsilon>0$ and let $\overline{S}_{n}=\overline{X}_{1,n}+\ldots+\overline{X}_{n,n}$. We note
$\mathbb{P}\!\left(\left|\frac{S_{n}-a_{n}}{b_{n}}\right|>\varepsilon\right)$ $\leq\mathbb{P}\big(S_{n}\neq\overline{S}_{n}\big)+\mathbb{P}\!\left(\left|\frac{\overline{S}_{n}-a_{n}}{b_{n}}\right|>\varepsilon\right).$
We bound the two terms separately. For the first term $\mathbb{P}(S_{n}\neq\overline{S}_{n})$, observe that the event $\{S_{n}\neq\overline{S}_{n}\}$ occurs if and only if at least one summand is truncated, that is, $\{S_{n}\neq\overline{S}_{n}\}=\bigcup_{k=1}^{n}\big\{|X_{n,k}|>b_{n}\big\}.$ Hence, by the union bound and assumption (i), $\mathbb{P}(S_{n}\neq\overline{S}_{n})\leq\sum_{k=1}^{n}\mathbb{P}\big(|X_{n,k}|>b_{n}\big)\xrightarrow{}0$ as $n\to\infty$.
For the second term, $\mathbb{P}\big(|(\overline{S}_{n}-a_{n})/b_{n}|>\varepsilon\big)$, note that $a_{n}=\mathbb{E}\,\overline{S}_{n}$. By Chebyshev’s inequality,
$\mathbb{P}\!\left(\left|\frac{\overline{S}_{n}-a_{n}}{b_{n}}\right|>\varepsilon\right)\leq\frac{1}{\varepsilon^{2}}\,\mathbb{E}\!\left[\left(\frac{\overline{S}_{n}-a_{n}}{b_{n}}\right)^{2}\right]=\frac{1}{\varepsilon^{2}b_{n}^{2}}\,\mathrm{Var}(\overline{S}_{n}).$
Because the variables $\overline{X}_{n,1},\dots,\overline{X}_{n,n}$ are independent and $\mathrm{Var}(Y)\leq\mathbb{E}Y^{2}$, we have $\mathrm{Var}(\overline{S}_{n})\leq\sum_{k=1}^{n}\mathbb{E}\,\overline{X}_{n,k}^{2}$. Combining these inequalities with by assumption (ii), we have
$\mathbb{P}\!\left(\left|\frac{\overline{S}_{n}-a_{n}}{b_{n}}\right|>\varepsilon\right)\leq\frac{1}{\varepsilon^{2}b_{n}^{2}}\sum_{k=1}^{n}\mathbb{E}\,\overline{X}_{n,k}^{2}\xrightarrow{}0,\text{ as }n\to\infty.\hskip 50.00008pt\blacksquare$
We can now apply this theorem to obtain the following result, also known as the Feller’s weak law of large numbers without a first moment assumption.
Theorem 10 (Durrett [2019] (Theorem 2.2.12))
Let $X_{1},X_{2},\ldots$ be i.i.d. with $x\operatorname{\mathbb{P}}(|X_{i}|>x)\to 0$ as $x\to\infty$. Let $S_{n}=X_{1}+\ldots+X_{n}$ and let $\mu_{n}=\operatorname{\mathbb{E}}X_{1}\boldsymbol{1}\{|X_{1}|\leq n\}$. Then $S_{n}/n-\mu_{n}\to 0$ in probability.
Proof: We apply Durrett’s Theorem 2.2.11 (i.e., Theorem 9) to the triangular array with entries $X_{n,k}=X_{k}$ for $1\leq k\leq n$ and the normalizing constants $b_{n}=n$. It suffices to check the two conditions of that theorem.
[…]
(ii)  Let $\overline{X}_{n,k}=X_{k}\mathbf{1}(|X_{k}|\leq n)$. We need to show $n^{-2}\sum_{k=1}^{n}\mathbb{E}(\overline{X}_{n,k}^{2})\;=\;n^{-1}\,\mathbb{E}(\overline{X}_{n,1}^{2})\ \longrightarrow\ 0$. By Lemma 2.2.13 of Durrett [2019] with $Y=|\overline{X}_{n,1}|$ and $p=2$,
$\mathbb{E}(\overline{X}_{n,1}^{2})=\int_{0}^{\infty}2y\,\mathbb{P}(|\overline{X}_{n,1}|>y)\,dy.$
Since $|\overline{X}_{n,1}|\leq n$, the integrand vanishes for $y\geq n$, and for $0\leq y\leq n$ we have
$\mathbb{P}(|\overline{X}_{n,1}|>y)\leq\mathbb{P}(|X_{1}|>y).$
Set $g(y)=2y\,\mathbb{P}(|X_{1}|>y)$. Hence
$\mathbb{E}(\overline{X}_{n,1}^{2})\leq\int_{0}^{n}g(y)\,dy.$
[…]
A.2 Weighted DKW Inequalities
With these results in place, we are ready to proceed with our proofs. Let $F_{Q}(x)=\operatorname{\mathbb{P}}_{Z\sim Q}(s(Z)\leq x)$ where $s(\cdot)$ denotes some fixed score function. For $Z_{i}\overset{\text{ i.i.d. }}{\sim}P$, we define
@@ A.2 Weighted DKW Inequalities
Assume that $Q$ is absolutely continuous with respect to $P$ ($Q\ll P$) and $dQ/dP\leq B$. Then, for all $\delta>0$
$\operatorname{\mathbb{P}}\Bigg(\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta+2C\sqrt{\frac{B}{n}}\Bigg)\leq 4e^{-n\delta^{2}/(2B^{2})},$
@@
$=\sup_{x\in\mathbb{R}}\left|\frac{1}{n}\sum_{i\in[n]}\frac{dQ}{dP}(Z_{i})\boldsymbol{1}\{s(Z_{i})\leq x\}-F_{Q}(x)\right|+\left|1-\frac{1}{n}\sum_{i\in[n]}\frac{dQ}{dP}(Z_{i})\right|.$ (17)
Regarding the first term, we have,
@@
$=\sup_{x\in\mathbb{R}}\left|\int f_{x}~d(P_{n}-P)\right|,$
@@
$1\leq\|dQ/dP\|_{P,2}^{2}=\int(dQ/dP)^{2}~dP\leq\|dQ/dP\|_{\infty}\int dQ/dP~dP\leq B.$
according to the Holder’s inequality. Hence, $\|dQ/dP\|_{P,2}\leq\sqrt{B}$ and by Theorem 7 we have
$\operatorname{\mathbb{E}}\left(\sup_{x\in\mathbb{R}}\left|\int f_{x}~d(P_{n}-P)\right|\right)\leq C\sqrt{\frac{B}{n}}\int_{0}^{1}\sqrt{1+\log N_{[]}\left(\epsilon\|dQ/dP\|_{P,2},\mathcal{F},L^{2}(P)\right)}~d\epsilon,$ (18)
with some universal constant $C>0$, and $N_{[]}\left(\epsilon\|dQ/dP\|_{P,2},\mathcal{F},L^{2}(P)\right)$ denotes the bracketting number for function class $\mathcal{F}=\{f_{x}:x\in\mathbb{R}\}$. It remains to compute an upper bound for $N_{[]}\left(\epsilon,\mathcal{F},L^{2}(P)\right)$. Let $t_{0}=$-$\infty$ and define,
$t_{i+1}:=\sup\left\{t\in\mathbb{R}:\int f^{2}(z)\boldsymbol{1}\{s(z)\in(t_{i},t]\}~dP<\epsilon^{2}\right\},$ (19)
@@
where $0\leq i\leq k-1$. Clearly, $\mathcal{F}\subset\cup_{i=0}^{k}b_{i}$ with $b_{k}=\{f\}$. Brackets $b_{0},\ldots,b_{k-1}$ have size
$\left(\int f^{2}(z)\boldsymbol{1}\{s(z)\in(t_{i},t_{i+1})\}~dP\right)^{1/2}\leq\epsilon,\quad 0\leq i\leq k-1,$
according to (19) and bracket $b_{k}$ has size 0. Hence, we have $k+1$ brackets in total with size smaller than $\epsilon$. By (20), we have $k-1\leq(\|dQ/dP\|_{P,2}/\epsilon)^{2}$ which implies that $N_{[]}\left(\epsilon,\mathcal{F},L^{2}(P)\right)\leq k+1\leq 2+(\|dQ/dP\|_{P,2}/\epsilon)^{2}$. Plugging this into (20), we get
@@
which follows from the observation that
@@
[…]
@@
Therefore, by McDiarmid’s inequality McDiarmid and others [1989] we get
@@
Back to (17), for the second term, Hoeffding’s inequality implies
@@
$\operatorname{\mathbb{P}}\Bigg(\sup_{x\in\mathbb{R}}\left|\hat{F}_{Q,n}(x)-F_{Q}(x)\right|>\delta\,\Bigg)\leq\frac{4\,CK}{\delta\sqrt{n}}+\frac{4(K^{2}-1)}{n\delta^{2}}.$
@@
Proof: By the same argument that led to (21) in the proof of Lemma 1, we get
@@
Proof: To prove this lemma, we compute the upper bound for the first term of (17) differently. According to Theorem 8, we get
@@
for $\epsilon\leq 2/3$, where $N_{[]}(\epsilon/8,\mathcal{F},L^{1}(P))$ denotes the bracketting number for function class $\mathcal{F}=\{f_{x}:x\in\mathbb{R}\}$. Similar to the proof of Lemma 1, it can be shown that $N_{[]}(\epsilon,\mathcal{F},L^{1}(P))\leq 2+1/\epsilon$. Therefore,
@@
Assume $Q\ll P$ and that ${dQ}/{dP}<\infty$ Q-a.s. Fix $B,\delta>0$ and $\epsilon\in(0,1/2)$. It holds that,
@@
$-\frac{8}{n\delta^{2}}{\operatorname{\mathbb{E}}\left(\left(\frac{dQ}{dP}(Z_{\text{calib}})\right)^{2}\boldsymbol{1}\left\{\frac{dQ}{dP}(Z_{\text{calib}})\leq n\right\}\right)},$
@@
$=\Bigg|$ $\operatorname{\mathbb{P}}\left(s(Z_{\text{test}})\leq x\Big|\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)\left(\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})\leq B\right)-1\right)$
@@
[…]
@@
$\leq\operatorname{\mathbb{P}}\left(\frac{dQ}{dP}(Z_{\text{test}})>B\right).$
@@
$\operatorname{\mathbb{P}}\bigg(\sup_{x\in\mathbb{R}}\left|F_{Q|\mathcal{X}_{B}}(x)-\hat{F}_{Q,n|\mathcal{X}_{B}}(x)\right|$ $\leq\delta+C^{\prime}\sqrt{B/n_{B}}\,\Big|\,n_{B}\bigg)\geq 1-4e^{-n_{B}\delta^{2}/(2B^{2})},$